GDP per capita in the United States was approximately $55,000 in 2015. Use the growth formula to answer the following questions: a. What will it be in the year 2020 if GDP per capita grows each year by 0 percent? GDP in 2020: b. What will it be in the year 2020 if GDP per capita grows each year by 2 percent? GDP in 2020: Growth Formula: (future value)-(present value)1 rMt present value this years GDP per capita future value GDP per capita in the future re rate of increase (in decimal form) per year t-number of years of growth
Answer #1

The formula for growth is F=P*(1+i)^n

If the growth rate is zero, then i=0,n is time which is 5 years and present GDP Per capita is 55000,so P=55000

So,

F=55000*(1+0)^5=55000

The GDP per capita growth will then be 55000 if the growth rate is zero. This is also obvious from the fact that growth rate is zero. That means the GDP will not grow but will be same in 2020 as it was in 2015.

If the growth rate is 2 percent

F=55000*(1+0.02)^5=60724.44